You are with your father with his motorbike on the way to your school, in the middle of your travel, the gas was empty. Luckily there is a gas station nearby. Your father asked you to buy 2 liters of 100 worth money per liters "premium" gasoline advertised with r=13. But you choose to buy 93 worth of money per 2 liters ordinary fuel with r=8. How much decrease in the ideal efficiency results from this decrease in the compression ratio? Use y=1.40. Is it greater that the difference in price?


1
Expert's answer
2021-04-08T09:02:23-0400

Thermal efficiency = "1 - (\dfrac{1}{r^{\gamma -1}})"


for r = 13

Thermal efficiency = "1 - (\dfrac{1}{13^{1.4-1}})" = "1 - (\dfrac{1}{13^{0.4}})" = 0.642


for r = 8

Thermal efficiency = "1 - (\dfrac{1}{8^{1.4-1}})" = "1 - (\dfrac{1}{8^{0.4}})" = 0.565


"\therefore" The decrease in efficiency is 0.642 - 0.565 = 0.077